Oscar Clemente is the manager of Forbes Division of Pitt, Inc., a manufacturer of biotech products. Forbes Division, which has $4 million in assets, manufactures a special testing device. At the beginning of the current year, Forbes invested $5 million in automated equipment for test machine assembly. The division’s expected income statement at the beginning of the year was as follows.

Sales revenue $ 16,000,000
Operating costs
Variable 2,000,000
Fixed (all cash) 7,500,000
Depreciation
New equipment 1,500,000
Other 1,250,000
Division operating profit $ 3,750,000

A sales representative from LSI Machine Company approached Oscar in October. LSI has for $6.5 million a new assembly machine that offers significant improvements over the equipment Oscar bought at the beginning of the year. The new equipment would expand division output by 10 percent while reducing cash fixed costs by 5 percent. It would be depreciated for accounting purposes over a three-year life. Depreciation would be net of the $500,000 salvage value of the new machine. The new equipment meets Pitt’s 12 percent cost of capital criterion. If Oscar purchases the new machine, it must be installed prior to the end of the year. For practical purposes, though, Oscar can ignore depreciation on the new machine because it will not go into operation until the start of the next year.

The old machine, which has no salvage value, must be disposed of to make room for the new machine.

Pitt has a performance evaluation and bonus plan based on residual income. Pitt uses a cost of capital of 12 percent in computing residual income. Income includes any losses on disposal of equipment. Investment is computed based on the end-of-year balance of assets, net book value. Ignore taxes.

Required:

a. What is Forbes Division’s residual income if Oscar does not acquire the new machine?

b. What is Forbes Division’s residual income this year if Oscar acquires the new machine?

c. If Oscar acquires the new machine and operates it according to specifications, what residual income is expected for next year?

(Enter your answers in thousands of dollars. Negative amounts should be indicated by a minus sign.)
Answer #1

Below Answer , please read very carefully . I already updated all formula in below worksheet . Any doubt , please let me know  

Oscar Clemente Amount ($)
Sales Revenue $    1,60,00,000
Operating Cost
Variable $       20,00,000
Fixed ( all cash)       75,00,000
Depreciation
New Equipment       15,00,000
Other       12,50,000
Div Operation Profit       37,50,000
Asset       40,00,000
Investment       50,00,000

ROI Calculation -

a)
What is Forbes Division’s residual income if Oscar does not acquire the new machine?
Operating Income /(Asset - Depreciation)+ ( Investment- Depreciation)
Amount ($)
Asset- a       40,00,000
Depreciation ( Others)-b       12,50,000
Net Balance ( a-b)=c       27,50,000
Investment (a)       50,00,000
Depreciation ( New Equipment )-b       15,00,000
Net Balance ( a-b)=d       35,00,000
Div Operation Profit       37,50,000
ROI 60%
ROI =Operating Income /(Asset - Depreciation)+ ( Investment- Depreciation)
3750000/(2750000+3500000)
B)
What is Forbes Division’s residual income this year if Oscar acquires the new machine?
Amount ($)
Asset- a       40,00,000
Depreciation ( Others)-b       12,50,000
Net Balance ( a-b)=c       27,50,000
Investment (d)       65,00,000
Net Investment (c+d)       92,50,000
Sale of Old Asset
Cost(e)       50,00,000
Depreciation ( Others)-f       15,00,000
Net Balance ( e-f)=g       35,00,000
Salvage Value                       -  
Loss on sale of Asset       35,00,000
Operating Profit       37,50,000
Loss on sale of Asset       35,00,000
( as above )
New Operating Income          2,50,000
Revised ROI
New Operating Income /Net Investment (c+d)
250000/9250000 2.70%

part C -

If Oscar acquires the new machine and operates it according to specifications, what residual income is expected for next year?
Revenue increase by 10%
Variable cost increased by 10%
Fixed cost decreased by 5%
Original $ Revised $
Sales Revenue $    1,60,00,000    1,76,00,000 (10% increased  
Operating Cost
Variable $       20,00,000       22,00,000 (10% increased  
Fixed ( all cash)       75,00,000       71,25,000 (5% decreased)  
Total cost       93,25,000
Amount($)
Depreciation
Current Asset       12,50,000
New
Cost       65,00,000
Salvage value 500000
Balance amount (a)       60,00,000
Depreciation (b) 3
Depreciation New per year(a/b)=C       20,00,000
Current Asset=D       12,50,000
Total Depreciation(C+D)       32,50,000
Revised Operating Margin
Sales Revenue $-A    1,76,00,000
Operating Cost
Total cost B( as above)       93,25,000
Total Depreciation C       32,50,000
( as above)
Revised Operating Margin=(A-B-C)       50,25,000
Investement Calculation
Asset ( a)       40,00,000 ( as above)
Depreciaiton ( 2 Year)(b)       25,00,000 ( as above)
( per year 1250000)
Net Balance (a-b)       15,00,000
New Investment (c)       65,00,000 ( as above)
Depreciaiton ( 1 Year)(d)       20,00,000 ( as above)
Net Balance (c-d)       45,00,000
Total Investment'       60,00,000
('1500000+4500000)
Revised ROI 84%
( Operating Profit/Net Investment)
('5025000/6000000)
( better to investment in New project with HIGHER Return )'
Amount ($)
Oprating Income          50,25,000 A
Depreciation ( as above)          20,00,000 B
Cost of New Machine
New Investment          65,00,000
Increase 12%
New Investment(a)          72,80,000
(6500000*112%)
Salvage value(b) 500000
Year 3
Depreciation (a-b)/ No of Year          22,60,000 C
Loss on Disposal of Old Machine          35,00,000 D
Revised Operating Income A+B-C-D
(A+B-C-D)= E          12,65,000
Old Machine Book$ Depreciation$ net Book Vlaue $
( as above )          40,00,000       25,00,000                 15,00,000
new Machine ( as above )          72,80,000       22,60,000                 50,20,000
Net Investment Value (F)                 65,20,000
ROI Operating Income / Net Investment
ROI E/F
(1265000/6520000) 19%